### v1

# SummExecEdit: A Factual Consistency Benchmark in Summarization with Executable Edits

Abstract
Detecting factual inconsistencies in summarization is critical, yet existing benchmarks lack the necessary challenge and interpretability for robust evaluation. In this paper, we introduce SummExecEdit, a novel benchmark leveraging executable edits to assess models on their ability to both detect factual errors and provide accurate explanations. The top-performing model, Claude3-Opus, achieves a joint detection and explanation score of only 0.49 in our benchmark, with individual scores of 0.67 for detection and 0.73 for explanation. Furthermore, we identify four primary types of explanation errors, with 45.4% of errors focusing on completely unrelated parts of the summary.
SummExecEdit: A Factual Consistency Benchmark in Summarization with Executable Edits
Onkar Thorat  Philippe Laban  Chien-Sheng Wu Salesforce AI Research {othorat, wu.jason}@salesforce.com
1 Introduction
Large language models (LLMs) have shown remarkable performance across various tasks by generating coherent responses while facing a major challenge with factual hallucination Huang et al. (2024). Various evaluation methods Laban et al. (2022); Fabbri et al. (2021, 2022); Tang et al. (2023); Luo et al. (2023); Yang et al. (2024) and inconsistency detection benchmarks have been proposed, and they used approaches such as editing ground truth texts to create intentional inconsistencies Laban et al. (2023) or generating benchmarks with LLMs where they label the factual accuracy of target texts post-generation. Tang et al. (2024b)
However, previous approaches usually have broad and sweeping edits, i.e., trivial edits and multiple edits in one sample. Kim et al. (2024) in their work chose to entirely rely on humans to generate edits, because initial experiments with LLM-based edits yielded mostly trivial edits. Such edits make factual inconsistency errors easier to be detected by LLMs, for example, GPT4 achieves 82.4% detection accuracy on a previous inconsistency benchmark - SummEdits Laban et al. (2023). This shows that we need mechanisms to generate more complex edits. Moreover, high quality explanations about inconsistency in these detection benchmarks are missing, and they are necessary to evaluate the model’s ability to explain their detection result and reason over facts.
To address these drawbacks, we propose to leverage the concept of executable edits Laban et al. (2024) to generate a challenging and interpretable benchmark called SummExecEdit - extending SummEdits Laban et al. (2023)  [^1] Our data is released at https://huggingface.co/datasets/Salesforce/summexecedit. While normal editing rewrites the entire summary with edits already applied, executable editing focuses on isolating and replacing a specific substring in the text showing what has changed, allowing precise changes that can introduce factual inconsistencies. Such edits helps LLM to better concentrate and come up with a complex yet granular and controlled edit with a meaningful explanation for same in a structured way, helping to eliminate broad and sweeping multiple edits, as shown in Figure 1.
We compare executable edits to existing editing and show that executable edits are superior, providing in range of 18-25% of higher number of challenging samples with better explanations. We evaluate a wide range of open-source and open-api access LLMs on SummExecEdit and find that most of these models lack on the combined task of detection and explanation of error, with the best model - Claude3-Opus achieving a joint score of only 0.49, with individual scores of 0.67 for detection and 0.73 for explanation. We also perform error analyses for incorrect explanations generated by candidate LLMs and categorize them into four reasons, with 45.4% of errors focusing on completely unrelated parts of the summary.
2 Related Works
Editing and Annotation in Previous Benchmarks. Several previous works have assessed various metrics in relation to human judgments associated with hallucination and factual inconsistencies. Works prior to advances in language models focused on human and expert level Fabbri et al. (2021); Falke et al. (2019) while the very recent works focused on generating annotations with the help of language models Chhikara et al. (2024); Tang et al. (2024b); Laban et al. (2023). Although these models generate coherent texts with a good level of inter-annotator agreements with respect to consistency Laban et al. (2023), such benchmarks may struggle to keep the changes in the original data to minimal Nguyen et al. (2024). Executable edits with help of language models was first explored by Laban et al. (2024), where it focused on replacing a substring for document editing. We build upon the works of SummEdits Laban et al. (2023) to use executable edits to generate new factual consistency benchmark.
Detecting Factual Inconsistencies. Since automatic evaluation metrics such as ROUGE Lin (2004) or BLEU Papineni et al. (2002) cannot reliably detect factual inconsistencies, several non-LLM based methods have been developed such as question-based Fabbri et al. (2022) and entailment-based Laban et al. (2022); Zha et al. (2023); Tang et al. (2024a). In various recent works, LLMs have shown superior performance when used as a judge to evaluate factual consistencies Luo et al. (2023). We explore this further in various settings and also re-evaluate the modern-day LLMs on our benchmark generated using executable edits.
3 Executable Edits
Given a (source, target) pair with a goal to edit the target such that it becomes factually inconsistent with the source, we can either ask the LLM to rewrite the target token-by-token, or ask it to isolate and highlight only a specific substring in the target text to be replaced. The latter can also be considered a program that can be executed and used. Executable editing is simple but it can help us generate more controlled and complex edits.
Figure 1: Example comparison between non-executable and executable edit.
Moreover, this minimizes synthetic data in the benchmarks since a majority of the original data remains same. Executable edits can also have a wide usage for data augmentation.
We first experiment and verify on a smaller sample to show that executable edits work better compared to non-executable edits. We selected around 100 original (document, summary) pairs from Laban et al. (2023) and used GPT4-Turbo, Claude3 Opus, and GPT3.5-Turbo to generate edited summaries and explanation of inconsistencies using both executable and non-executable prompts (Appendix F) in a structured json format. This generates around 600 edits, which are shuffled, anonymized, and annotated by two of the authors manually based on the four questions: a) is the edit inconsistent? b) is the edit complex/good quality? c) is the edit controlled/granular? d) is the explanation quality good? Each annotator annotated nearly 400 edits with 200 in common to verify the inter-annotator agreement in Table 3.
Condition N %Controlled ($\uparrow$) %Inconsistent ($\uparrow$) %Complex ($\uparrow$) %Explanation ($\uparrow$)
GPT4-Turbo (Exec) 144 86.81 81.25 45.83 44.44
GPT4-Turbo (Non-Exec) 134 90.30 81.34 23.13 20.90
GPT3.5-Turbo (Exec) 134 76.87 73.13 17.16 16.42
GPT3.5-Turbo (Non-Exec) 133 86.47 72.18 18.05 12.03
Claude3-Opus (Exec) 138 92.03 84.78 49.28 48.55
Claude3-Opus (Non-Exec) 136 97.06 88.24 31.62 30.15
Table 1: Manually annotated scores for comparison between executable and non-executable edits. Each column shows percentage of N but does not consider the edits filtered out in its previous columns.
The result for our manual annotation is given in Table 1. We use a filtering mechanism - each subsequent column filters out the edits deemed inappropriate by either of the annotators in the previous column. The models show a similar trend - executable edits lead to a higher score towards the end implying a higher number of good edits and explanations. For example, Claude3-Opus provides nearly 18% more controlled and high quality edits with executable prompt.
4 SummExecEdit Benchmark
We leverage Laban et al. (2023) to build our benchmark across 10 domains such as News, Podcast, Bill, Sales calls, etc. Based on Table 1, we use both GPT4-Turbo and Claude3-Opus models to generate our benchmark using the executable prompt. Both the models are asked to generate six edits for each (document, summary) pair.
After generating these edits, we remove the trivial edits with help of GPT4-Turbo. We ask GPT4-Turbo to classify an edit as a date change, number change, antonym change, or others. Any edit classified as date, number, or antonym change is removed from the benchmark.
The final benchmark results in 2,121 factually inconsistent summaries. To balance out the entire benchmark, we use all 2,120 factually consistent edits from SummEdits benchmark resulting in a total of 4,241 samples in the final benchmark. Each of the 10 domains provide around 200-300 inconsistent summaries. The distribution of each domains is given in Table 6.
5 Results
To evaluate LLMs on SummExecEdit benchmark, we use two types of prompts. D&E - Detect and Explain error. Models need to detect if there is any factual inconsistency in summary, if yes, explain the inconsistency (Appendix F.3). E|D - Explain error given Detection. Given that the summary is inconsistent, models need to explain the inconsistency in the summary (Appendix F.4).
Table 2 provides the Detection Accuracy (DA) of all the models using prompt D&E. The best performing model on SummExecEdit , GPT4o, provides an accuracy of only around 73%. The overall results show that many LLMs struggle in detecting the factual error. As a reference, we also evaluate two non-LLM based approaches which use far lesser compute - AlignScore Zha et al. (2023) receiving 57.4% and MiniCheck Tang et al. (2024a) receiving 60.0% accuracy, which are better than several open-source LLMs.
5.1 Evaluating Explanations
Two of the authors manually annotated around 1200 explanations with 300 in common. The results for manual annotation are given in appendix in Table 5. Based on these annotations we evaluated different LLM-as-Judge Zheng et al. (2023) by asking them to assign a label as Entirely Correct (1), Partially Correct (0.5), or Not Correct (0). We use following four types of prompts for explanation evaluation using GPT4o, GPT3.5-Turbo, Claude3-Opus, and Claude3-Haiku.
  • Reference less and no seed (EvalV1) - We provide document, edited summary, and LLM’s explanation to be evaluated.
  • Reference less and with seed (EvalV2) - We provide seed summary (ground truth), edited summary, and LLM’s explanation to be evaluated.
  • Reference based with seed (EvalV3) - We provide seed summary (ground truth), edited summary, reference explanation, and LLM’s explanation to be evaluated.
  • Reference only (EvalV4) - We provide reference explanation and LLM’s explanation to be evaluated.
The correlations for four prompts with different LLMs with respect to our manually annotated explanations are provided in appendix in Table 4. Selected explanations for manual annotations were shuffled and randomly selected. The model and either of the two prompts that generated those explanations were anonymized. The IAA between both the annotators are - Correlation of 0.885 and Cohen Kappa of 0.81. We think that the reference explanations generated at the time of edit itself are the best and of high quality, scoring 0.95 for 40 samples. The prompt EvalV4 works the best which suggests that evaluating explanations or reasoning of models, works better when we have access to what the edit/reason is or a reference explanation.
We use prompt EvalV4 with GPT4o model to give the Explanation Score (ES) of all models on 2,121 inconsistent summaries generated using both our prompts - D&E and E|D. The results are provided in Table 2.
D&E E|D
Model DA($\uparrow$) DS($\uparrow$) ES($\uparrow$) JS($\uparrow$) ES($\uparrow$)
claude3-opus 0.71 0.67 0.733 0.491 0.684
gemini-1.5-pro 0.728 0.728 0.665 0.484 0.628
gpt-4o 0.733 0.725 0.629 0.456 0.684
gpt4-turbo 0.729 0.581 0.714 0.415 0.711
gemini-1.5-flash 0.708 0.556 0.633 0.352 0.62
gpt3.5-turbo 0.585 0.676 0.467 0.315 0.488
command-r-plus 0.655 0.369 0.703 0.259 0.595
gemini-pro 0.625 0.432 0.499 0.216 0.456
claude3-haiku 0.622 0.54 0.387 0.209 0.473
claude3-sonnet 0.636 0.33 0.626 0.206 0.633
command-r 0.609 0.414 0.489 0.202 0.49
palm-bison 0.608 0.248 0.473 0.117 0.407
mixtral-8x7b 0.609 0.73 0.503 0.367 0.568
mistral-large 0.703 0.512 0.714 0.366 0.624
llama3-70b 0.678 0.423 0.713 0.302 0.585
llama3-8b 0.555 0.73 0.239 0.175 0.458
mistral-7b 0.541 0.107 0.735 0.079 0.488
Table 2: Scores for D&E (Prompt Detect and Explain error) and E|D (Prompt Explain error given Detection) for inconsistent summaries. DA - Detection Accuracy, DS - Detection Score, ES - Explanation Score, JS - Joint Score. ES evaluated by EvalV4 prompt using GPT4o. DA based on entire benchmark. DS, ES, and JS based only on inconsistent summaries. Last 5 models are open-source.
5.2 Joint Scores on SummExecEdit
We define the joint scores comprising of both detection and explanation scores on factually inconsistent summaries in SummExecEdit . Detection Score (DS) is calculated only on 2,121 factually inconsistent summaries, and we give a score of 1 if model correctly detects the summary being factually incorrect and 0 otherwise. Joint score (JS) is calculated by multiplying both DS and ES element-wise, and the results are presented in Table 2.
The best model Claude3-Opus achieves a JS of 0.49 which suggests that the task of detecting factual inconsistency and explaining the same is still a challenging task for most modern-day LLMs, making them incapable to reason out-of-the-box. It is worth noting the big JS gap for open-api and open-source models. At the same time, it is also good to see Mixtral-8x7b achieving the best DS. This also brings up an interesting finding - some models are good at detecting the factual errors but struggle to explain the error, and also vice-versa. Models belonging to the same family also show differing behavior, for example GPT-4o and GPT4-Turbo or Mixtral-8x7b and Mistral-Large show different trends.
5.3 Error Analysis for Explanations
We analysed 350 of our manually annotated explanations that were incorrect or partially correct and observed that most of the errors in these explanations mainly fall under the following categories. If an explanation contains multiple errors, we report the first found error in the explanation.
Misattribution of Error - This is the most common type of error, accounting for 45.4% of incorrect explanations. The explanation would focus on a completely unrelated part of the summary or the document and assign the blame on it.
Additional Unrelevant Explanation - The LLM provides the correct explanation but also continues to generate some unrelated explanation. Such explanations make 28.9% of incorrect explanations.
Concentrating on Completeness - The explanation focuses on completeness showing missing details in summary rather than focusing on factual correctness. This accounts for 15.4% of incorrect explanations.
Vague Explanation - These are either complex to understand or incomplete explanations missing out on details. They may correctly identify the error but not effectively explain it. They account for 10.3% of incorrect explanation.
While we do not find any relation of errors with specific models or specific prompts, we find a relation where different models happen to make similar errors in explanations belonging to the same document and factually inconsistent summary pairs.
6 Conclusion
In this work, we first explored the executable editing with LLMs for generating benchmarks. Through our experiments, we show its superiority in generating a challenging benchmark and use it to generate a new factual consistency detection and explanation benchmark - SummExecEdit . We evaluate various LLMs over this benchmark for both detection and explanation of factual inconsistencies. The best performing model for detection - GPT4o achieves an accuracy of 73% while Claude3-Opus performs best on combined task of detection and explanation achieving a joint score of 0.49. These scores suggest that SummExecEdit is a challenging benchmark. We explore various prompt formats that lead to better auto-evaluation by LLMs and perform error-analysis for incorrect explanations generated by LLMs. We hope researchers and practitioners use executable edits to generate and augment data, and SummExecEdit can be used by LLM Developers for evaluating LLMs’ abilities to detect factual errors and reason about facts.
Limitations
While executable edits help in creating more challenging benchmark, it is important to note a few limitations of these edits and this work in general. First, generating benchmarks with executable edits require availability of ground truth (source, target) pairs which might not always be the case. Second, non-executable edits are equally capable of generating controlled edits but these are simple and low quality. If the end-goal is to only generate controlled edits and complexity does not matter, non-executable edits could be used as well. Last, we have only experimented using these edits for creating a single benchmark related to summarization. Further experiments in other domains, as well as for data augmentations, are required.
Appendix A Model Details
OpenAI Models. We used three OpenAI models in our experiments - GPT4o, GPT4-Turbo, and GPT3.5-Turbo. All the models were accessed from the official OpenAI API.
Google Models. We used four Google models in our experiments - Gemini-1.5-Pro, Gemini-1.5-Flash, Gemini-Pro, and Palm-Bison. We used Google Cloud Platform to access Google models, which relies on the VertexAI API.
Anthropic, Cohere, Llama3, Mistral models. We used the Claude models in Anthropic family, Command-R models in Cohere family, models in Llama3 family, and models in Mistral family on AWS bedrock.
Appendix B License Details and Terms of Use
We leverage Laban et al. (2023) for raw data which was released under Apache-2.0 license, granting copyright license.
Appendix C Annotation Tool
Figure 2 provides an example of our annotation tool. The text highlighted in red, is replaced by the text highlighted in green and on the right side we perform the required annotation. This helps us annotate more reliably and easily.
Figure 2: Two-Column annotation interface with highlights used for annotation
Appendix D Manual Annotation and Inter-Annotator Agreement
In Section 3, we manually annotate the generated inconsistent summaries to compare between executable and non-executable edits based on following 4 questions:
Is the edit inconsistent? - We annotate if the edited summary is factually inconsistent or the replaced text still kept the edited summary consistent.
Is the edit complex/good quality? - We annotate if the factual inconsistency is complex or interesting, meaning the edits should not be trivial such as easy antonym swap, negating sentence, or numerical and date changes. Trivial edits make the benchmark easier.
Is the edit controlled/granular? - We annotate if the edit is controlled and granular or still very broad and sweeping, a drawback of previous benchmarks.
Is the explanation quality good? - We annotate if the reference explanation specifying the reason of inconsistency generated by LLM at the time of edit is correct.
Table 3 provides the Cohen Kappa as the inter-annotator agreement.
Question N Cohen Kappa
Inconsistent 202 0.76
Interesting 177 0.61
Controlled 156 0.6
Explanation quality 43 0.49
Table 3: Inter-annotator agreement between both the authors. We keep filtering out the examples in every row to remove the bias of non-matching annotations since the subsequent annotation of questions for every example depend on the previous questions. The Cohen for Explanation looks low but the disagreement is on a very small samples. Since Cohen Kappa is looking at classes with equal weights and due to a big imbalance, it has a larger effect.
Table 4 provides the Correlation and Balanced Accuracy for manual annotation of explanations based on EvalV1, EvalV2, EvalV3, and EvalV4 for four models under consideration.
Correlation BAcc
Model V1 V2 V3 V4 V1 V2 V3 V4
claude3-haiku 0.076 0.478 0.732 0.725 0.385 0.49 0.655 0.618
claude3-opus 0.484 0.782 0.801 0.774 0.46 0.76 0.776 0.738
gpt3.5-turbo 0.103 0.118 0.325 0.563 0.375 0.401 0.495 0.625
gpt4o 0.603 0.816 0.804 0.833 0.533 0.792 0.794 0.811
Table 4: Correlation and Balanced Accuracy of prompts and models for explanation evaluation with respect to manual evaluation.
Table 5 provides explanation scores for our manual annotations of explanation evaluation. We can see a comparable performance gap between the open-source and open-api access models.
Prompt V1 Prompt V2
Model N ES($\uparrow$) N ES($\uparrow$)
claude3-opus 29 0.897 39 0.884
gpt-4o 28 0.821 39 0.884
gpt4-turbo 26 0.807 20 0.8
mixtral-8x7b 31 0.516 40 0.662
gpt3.5-turbo 30 0.567 40 0.7
llama3-70b 20 0.775 40 0.725
command-r-plus 17 0.706 40 0.762
gemini-pro 17 0.676 40 0.587
claude3-haiku 25 0.54 40 0.537
claude3-sonnet 19 0.763 40 0.725
command-r 15 0.6 40 0.625
llama3-8b 36 0.319 40 0.625
palm-bison 10 0.6 40 0.6
mistral-7b 6 0.833 40 0.612
Table 5: Scores for prompt V1 and V2 for inconsistent summaries. N - Number of samples, ES - Explanation Score. ES for Manual Annotation based on 887 explanations annotated by one author. The score for 40 manually annotated reference explanation is 0.95.
Appendix E SummExecEdit Statistics
Table 6 provides statistics for different domains present in the DEFES benchmark.
Domain N %Inconsistent
SciTLDR 307 58.31
News 567 47.8
Podcast 344 58.72
BillSum 608 44.9
SamSum 450 51.11
Shakespeare 511 41.3
QMSum 334 50.3
ECTSum 416 47.12
Sales Email 368 57.1
Sales Call 336 53.87
Total 4241 50
Table 6: Statistics of the domains in SummExecEdit
Appendix F Prompts
F.1 Executable Prompt
You are given a document and its corresponding summary. You must generate 6 edits to the summary which should strictly follow the guidelines mentioned below.
While making edits, think of what in a summary could be changed, added, or removed that is difficult to detect but makes the summary inconsistent with the document. Be mindful, that you cannot change the figures, numbers, dates, digits mentioned in the summary.
The edit should introduce an error which makes the summary factually incorrect, or inconsistent with the document. Few examples of good error types are given after the guidelines. Refer to those edits and generate similar error types. For each edit only one of the error type should be used.
[Guidelines]
- You can generate only one error per edited you generate.
- Your output should be a valid JSON string: that starts with : {"edits": [ ... ]} and contains edits in the format described below.
- Remember to not introduce following types of errors:
- You cannot use antonyms or opposite words to introduce an error.
- You cannot negate the sentence to introduce an error.
- You cannot replace a date by some other random date or random time of the year to introduce an error.
- You cannot replace some amount or money by other random amount or money.
- You cannot replace a number or a percentage by some other random number or percentage to introduce an error.
- You cannot replace a name by some other random name to introduce an error.
- Your explanation should not mention the original summary or the edit, it should only describe the difference between the new summary and document with respect to the document: "The document says X but the document says Y."
- Make sure your "original_text" is a substring of the summary.
- You should diversify the types of errors you generate as much as possible. Do not generate all edits of the same type.
- In the explanations you generated, you must refer to the edited summary simply as the summary. You must not mention edits or editing, or the "edited summary" in any way, simply call it the summary.
[Edit format]
Each edit should be a valid JSON object, with three keys as show: "original_text", "replace_text", and "explanation". The "original_text" key should be the original text in the summary that you are editing which is a substring of the summary. The "replace_text" key should be the new text that you are replacing the original text with. The "explanation" key should be a 1-2 sentence explanation of why the edit is an error.
[Document]
[DOCUMENT]
[Summary]
[SUMMARY]
F.2 Non-Executable/Normal Prompt
You are given a document and its corresponding summary. You must generate 6 edits to the summary which should strictly follow the guidelines mentioned below.
While making edits, think of what in a summary could be changed, added, or removed that is difficult to detect but makes the summary inconsistent with the document. Be mindful, that you cannot change the figures, numbers, dates, digits mentioned in the summary.
The edit should introduce an error that makes the summary factually incorrect, or inconsistent with the document.
[Guidelines]
- You can generate only one error per edited you generate.
- You must not rewrite an entirely new summary. Your edited summary should modify, insert or delete AT MOST 5 words of the original summary.
- Remember to not introduce following types of errors:
- You cannot use antonyms or opposite words to introduce an error.
- You cannot negate the sentence to introduce an error.
- You cannot replace a date by some other random date or random time of the year to introduce an error.
- You cannot replace some amount or money by other random amount or money.
- You cannot replace a number or a percentage by some other random number or percentage to introduce an error.
- You cannot replace a name by some other random name to introduce an error.
- You should diversify the types of errors you generate as much as possible. Do not generate all edits of the same type.
- You do not need to mention the type of edit that you made.
- For each of the three edits, rewrite the entire summary with the edit you make.
- In the explanations you generated, you must refer to the edited summary simply as the summary. You must not mention edits or editing, or the "edited summary" in any way, simply call it the summary.
[Edit format]
You should generate six edits to the original summary, and for each generate an one-sentence explanation of what the error in the edited summary is.
Each edit should follow following json format:
{"edits": [
{"edited_summary": "[Edited version of the summary]", "explanation": "A natural language explanation of the error."},
[Document]
[DOCUMENT]
[Summary]
[SUMMARY]
F.3 Detect and Explain error (D&E)
You are given a document and its corresponding summary which may or may not be factually correct and consistent with the document. Your task is to generate a valid json string (use escape characters if double quotes and new lines are used within value) that has following two fields:
1. "consistent" - This field gives whether the summary is factually correct and consistent with respect to the document. It should be "yes" if the summary is factually consistent with respect to the document and "no" otherwise.
2. "explanation" - If you set "consistent" to "no", then this field should give the explanation as why you think it is incorrect or inconsistent. If you set "consistent" to "yes", then the explanation should be an empty string. Example format:
{"consistent": "no", "explanation": "[...]"}
or
{"consistent": "yes", "explanation": ""}
Now complete the task for the following document, summary pair:
[Document]
[DOCUMENT]
[Summary]
[SUMMARY]
F.4 Explain error given Detection (E|D)
You are given a document and its corresponding summary. We know that the summary contains a factual error which renders it inconsistent with the facts in the document. Your task is to provide a 1-2 sentence explanation that identifies what fact or facts in the summary is/are inconsistent with the document. You should output your explanation in a valid json format (use escape characters for double quotes and new lines if used within value), in the following format:
{"explanation": "[...]"}
Now complete the task for the following document, summary pair:
[Document]
[DOCUMENT]
[Summary]
[SUMMARY]